if $ 10,000 is invested in an account at 3% annual interest compounded quaterly, how much will be in the account in 5 years if no money is withdrawn?
Answer by lwsshak3(11628)

if $ 10,000 is invested in an account at 3% annual interest compounded quaterly, how much will be in the account in 5 years if no money is withdrawn?
***
compound Interest formula: A=P(1+i)^n, i=interest rate per period, P=initial investment, n=number of periods, A=amount after n periods.
For given problem:
P=$10,000
i=.03/4=0.0075
n=5*4=20
..
A=10000(1+.0075)^20
A=10000(1.0075)^20
A=10000*1.1612
A=11,612
how much will be in the account in 5 years=$11,612